OpenStudy (anonymous):

sin^2 theta-cos^2 theta=0

OpenStudy (vishweshshrimali5):

\[\large{\sin^2{\theta} - \cos^2 {\theta} = 0}\] I can write this as : \[\large{\cos^2 {\theta} - \sin^2{\theta} = 0}\]

OpenStudy (vishweshshrimali5):

Then, I use this formula: \[\large{\cos (2x) = \cos^2 x - \sin^2 x}\] Thus, I get, \[\large{\cos(2\theta) = 0}\]